OpenStudy (anonymous):

Calculate the following integrals \[\Large \int_0^{\frac{\pi}{2}} dx \cos(x)e^{-sin(x)}\]

OpenStudy (anonymous):

let u=sin x

OpenStudy (australopithecus):

use substitution u = sin(x) du = cos(x)dx

OpenStudy (australopithecus):

you can just sub du in and eliminate the cos(x)dx

OpenStudy (anonymous):

guys i am quite very new to integrals i need complete solution rather then advice, i have exam tomorrow this question possible to appear on my exam paper

OpenStudy (australopithecus):

crap

OpenStudy (australopithecus):

\[\int\limits_{0}^{\pi/2}e^{-u}du\]

OpenStudy (australopithecus):

can you compute the antiderivative?

OpenStudy (australopithecus):

anti derivative of e^(-u) = -e^(-u)

OpenStudy (australopithecus):

then resub in u=sin(x) you will get this after replacing u